# HealthQ: Unveiling Questioning Capabilities of LLM Chains in Healthcare Conversations

### v3

@@ Abstract
In digital healthcare, large language models (LLMs) have primarily been utilized to enhance question-answering capabilities and improve patient interactions. However, effective patient care necessitates LLM chains that can actively gather information by posing relevant questions. This paper presents HealthQ, a novel framework designed to evaluate the questioning capabilities of LLM healthcare chains. We implemented several LLM chains, including Retrieval-Augmented Generation (RAG), Chain of Thought (CoT), and reflective chains, and introduced an LLM judge to assess the relevance and informativeness of the generated questions. To validate HealthQ, we employed traditional Natural Language Processing (NLP) metrics such as Recall-Oriented Understudy for Gisting Evaluation (ROUGE) and Named Entity Recognition (NER)-based set comparison, and constructed two custom datasets from public medical note datasets, ChatDoctor and MTS-Dialog. Our contributions are threefold: we provide the first comprehensive study on the questioning capabilities of LLMs in healthcare conversations, develop a novel dataset generation pipeline, and propose a detailed evaluation methodology.
keywords:
Artificial Intelligence in Healthcare , Biomedical Informatics , Digital Health , Natural Language Processing , Information Retrieval in Medicine , Medical Conversational AI , Healthcare Informatics
[…]
[UCI]organization=University of California, Irvine, addressline=Department of Computer Science, city=Irvine, postcode=92697, state=CA, country=USA
[…]
[UCI2]organization=University of California, Irvine, addressline=School of Nursing, city=Irvine, postcode=92697, state=CA, country=USA
[…]
[Columbia]organization=Columbia University, addressline=Department of Statistics, city=New York, postcode=10027, state=NY, country=USA
[…]
[WUSTL]organization=Washington University in St. Louis, addressline=Department of Computer Science and Engineering, city=St. Louis, postcode=63130, state=MO, country=USA
1 Introduction
@@
The integration of Artificial Intelligence (AI) in healthcare has significantly transformed the landscape of personal healthcare, enabling advancements in diagnostic accuracy, treatment personalization, and patient management [1, 2, 3, 4]. Central to these advancements are Large Language Models (LLMs), which have shown outstanding abilities in comprehending and producing human-like text. These capabilities have been harnessed to facilitate various healthcare interventions, from virtual health assistants to automated diagnostic systems [5]. Current research in digital healthcare has extensively explored the use of LLMs to enhance question-answering (QA) capabilities. QA systems are crucial as they provide patients with accurate and timely responses to their medical inquiries, thereby improving patient outcomes, increasing accessibility to healthcare resources, and enhancing overall patient satisfaction and well-being. Numerous datasets and benchmarks, such as MedQA [4], emrQA [6], and MedMCQA [7], have been developed to train and evaluate LLMs in their ability to answer patient questions effectively.
Recent advancements demonstrate LLMs can be organized into chains to tackle more complex interactions [5], for instance, interacting with patients via natural language while utilizing components that enhance the capabilities of LLMs by integrating additional tools and knowledge bases, such as Retrieval-Augmented Generation (RAG) [8] and ReAct [9]. This structure allows the LLM to not only generate answers to questions but also retrieve relevant information, apply reasoning, and perform specific tasks more effectively.
In real-world LLM healthcare applications, the capability to answer patient questions is just one facet of their functionality [10]. Effective interaction requires LLM healthcare chains to actively engage patients by asking relevant and context-aware questions to gather comprehensive information. This includes symptoms, medical history, lifestyle factors, and other pertinent details that are critical for accurate diagnosis and personalized treatment. By actively soliciting detailed patient information, LLM healthcare chains can enrich the context in which they operate, leveraging integrated tools and smart agents to enhance diagnostic accuracy, provide better interventions, and offer more tailored health advice. Despite the critical importance of this interactive capability, the evaluation of LLM healthcare chains’ questioning abilities—specifically, their effectiveness in asking pertinent questions to elicit essential information from patients—remains significantly under-explored [11, 12].
Addressing this gap, our work proposes a novel evaluation framework to assess and benchmark the questioning capabilities of LLM healthcare chains during conversational interactions with patients. HealthQ not only evaluates the quality of the questions posed by the LLM healthcare chains but also examines whether these questions help in obtaining better answers. We employ mutual information validation to demonstrate the relationship between the quality of the questions and the improvement in the answers received. This approach ensures that the LLM healthcare chains can effectively interact with patients, leading to more accurate diagnoses and better healthcare outcomes.
In summary, our contributions are threefold:
@@
    We conduct the first comprehensive study on the questioning capabilities of LLM healthcare chains, filling a significant gap in healthcare AI research.
@@
    We introduce a novel dataset generation pipeline, specifically designed to evaluate the question-asking performance of LLM healthcare chains in realistic patient interaction scenarios.
@@
    We develop and rigorously validate an evaluation framework that integrates multiple LLM chain types and traditional Natural Language Processing (NLP) metrics. Our framework is validated through two key research questions, ensuring its robustness and providing a solid benchmark for future studies on LLM question-asking in healthcare.
We aim for this framework to serve as a platform for researchers to develop general chain models capable of solving complex real-world questions through extensive logical reasoning in both retrieval and comprehension stages. This approach will enhance the effectiveness and reliability of LLM healthcare chains in diverse healthcare scenarios. To further contribute to this area, we will publish our code and datasets.
2 Related Work
2.1 Digital Healthcare
The advent of digital healthcare interventions has revolutionized patient care by leveraging technologies such as wearable devices [13, 14, 15], AI-driven analytics [5, 16], and advanced decision-making frameworks for healthcare resource allocation and optimization [17, 18, 19, 20]. These interventions enable remote monitoring, personalized treatment plans, and predictive analytics, improving patient outcomes and reducing healthcare costs [21, 22]. AI integration, including machine learning and natural language processing for diagnostic accuracy and patient management [23, 1], as well as parameter-efficient fine-tuning approaches [24], are further enhancing healthcare efficiency. Moreover, modeling techniques and representation learning have shown promise in optimizing healthcare systems [3], emphasizing the transformative potential of digital healthcare.
2.2 LLMs in Healthcare
LLMs such as GPT-4 have shown remarkable proficiency in generating human-like text and understanding complex language constructs [25]. These capabilities have been harnessed in various healthcare applications, including automated medical record summarization, virtual health assistants, and clinical decision support systems [26, 27]. Automated medical record summarization condenses patient records into concise summaries, saving time for healthcare providers. Virtual health assistants interact with patients using natural language, providing answers to medical queries and reminders for medication, thereby enhancing patient engagement. Clinical decision support systems leverage the extensive knowledge embedded in LLMs to assist in diagnosing diseases and suggesting treatment options. LLMs streamline clinical workflows and improve patient engagement by providing accurate and timely responses to medical queries, ultimately enhancing the overall quality of care.
2.3 Interactive Questioning with LLM Healthcare Chains
@@
Several approaches have been introduced to refine the questioning capabilities of LLMs. Techniques such as Chain of Thought (CoT) [29] prompting, reflective chains [30], and Retrieval-Augmented Generation (RAG) [8] integrate reasoning processes that help generate more contextually relevant questions. CoT prompting allows LLMs to break down complex queries into intermediate steps, leading to more precise and targeted questioning. Reflective chains enable the model to iteratively improve the questions by reflecting on previous intermediate answers, thus enhancing the relevance of subsequent queries. RAG, on the other hand, combines retrieval and generation processes to ensure that questions are informed by the most relevant information retrieved from a knowledge base. Despite these advancements, most existing work has primarily focused on optimizing task completion and response accuracy, with less attention given to the development of sophisticated questioning strategies that can drive more informative patient interactions. Our work addresses this gap by specifically evaluating the questioning capabilities of LLM healthcare chains, contributing a novel framework that benchmarks their ability to engage in meaningful and effective patient dialogues.
3 Method
3.1 Overview of HealthQ
@@
3.2 Data
HealthQ necessitates a general data processing mechanism due to the unstructured nature of medical notes and the often ambiguous doctor-patient conversations. Patients may provide simple ”yes” or ”no” answers, leaving the context or symptoms embedded in the doctor’s questions. To accurately evaluate the questioning capabilities of LLM healthcare chains, we require a customized data format that aligns with our evaluation framework. This involves clearly separating patient known knowledge—which includes the patient’s feelings, self-reported symptoms, and background history—from the doctor’s question statements. These pairs are essential for our virtual patient mechanism, enabling the assessment of how effectively doctors ask questions to gather useful information for diagnosis. To illustrate how the dataset format changed through our processing, Table 1 shows an example of the original and processed data formats.
3.2.1 Dataset
@@
Table 1: Example of original and processed data. The original format includes detailed patient background, symptoms, diagnosis, history, plan of action, and dialogue. The processed format extracts essential patient known knowledge and reformulates doctor question statements for clarity and consistency.
@@
3.2.2 Data Processing
As depicted in Figure 1, we split the original medical notes into training and test sets. The training set is transformed and stored in a vector database (VDB), which the LLM Chain can use for searching and retrieval. However, the LLM Chain has no knowledge of any new patients in the test set. The test set is parsed by the LLM to simulate virtual patients, allowing us to evaluate how effectively these LLM healthcare chains, acting as doctors, can ask relevant questions and make judgments based on the simulated patient responses.
In more specific, each training dataset is processed to create a VDB for searching, implemented using FAISS [^4] https://github.com/facebookresearch/faiss. The customized collate functions can specify comparison fields and utilize metadata for enhanced information retrieval. This flexibility accommodates diverse datasets with varying forms and topics. During agent executions, multiple VDB searching functions can be provided as tools without merging, allowing the LLM to select the appropriate database for searching, also allowing for heterogeneous search functions. VDBs work by finding the nearest vectors, which are represented by embedding models. Each vector database entry is defined as:
@@ 3.2.2 Data Processing
Each entry in the vector database, referred to as a record, is structured to include both the content and the associated metadata. The collate function processes the training data to create these records:
$\text{Record}_{i}=(\text{Content}_{i},\text{Metadata}_{i})=\text{Collate}(\text{Data}_{\text{train}}(i))$
These databases were critical in facilitating efficient retrieval and processing of medical information, ensuring that the LLM health chains had access to relevant and structured data during their training and evaluation phases. To generate ground truth data from the test set, we employed the Claude [^5] https://claude.ai/ to parse patient notes into structured formats. The parsing function processes the input patient data to produce an output pair consisting of patient known knowledge and doctor question statements.
The final ground truth for testing, $G$, is thus composed of the results from this processing pipeline:
$GT=\{\text{LLM\_Parser}(\text{Data}_{\text{test}}(i))\}$
3.3 Baseline LLM Healthcare Chains
HealthQ includes several state-of-the-art chains, each with unique mechanisms for generating diagnostic questions. Below are the detailed descriptions of each chain implementation, followed by a combined pseudo-code representation.
@@ 3.3 Baseline LLM Healthcare Chains
Table 2: LLM Judge Interrogation Metrics (Scores range from 0 to 10)
Metrics Description
Specificity Measures how specifically the question targets the patient’s symptoms and conditions.
Usefulness Assesses the utility of the question in providing relevant information for diagnosis.
Relevance Evaluates the pertinence of the question to the patient’s symptoms.
Coverage Determines the extent to which the question covers the information available in the ground truth data.
Fluency Rates how well the question is formulated in fluent, natural language.
3.4 Evaluation Framework
@@
  • 2.
    LLM Judge Interrogation: This metric evaluates the questions generated by the LLM healthcare chains based on specificity, usefulness, relevance, coverage, and fluency, as shown in Table 2. An LLM judge $J$ assesses each question $Q$ using the patient knowledge $K$, providing scores according to these criteria. The interrogation metrics are calculated as follows:
@@
    Evaluate the Question ($Q$) using Judge ($J(Q,K)$): Evaluate the generated question using an LLM judge based on criteria such as relevance, specificity, and informativeness, as detailed in Table 2.

### v4

@@ Abstract
Effective patient care in digital healthcare requires large language models (LLMs) that not only answer questions but also actively gather critical information through well-crafted inquiries. This paper introduces HealthQ, a novel framework for evaluating the questioning capabilities of LLM healthcare chains. By implementing advanced LLM chains, including Retrieval-Augmented Generation (RAG), Chain of Thought (CoT), and reflective chains, HealthQ assesses how effectively these chains elicit comprehensive and relevant patient information. To achieve this, we integrate an LLM judge to evaluate generated questions across metrics such as specificity, relevance, and usefulness, while aligning these evaluations with traditional Natural Language Processing (NLP) metrics like ROUGE and Named Entity Recognition (NER)-based set comparisons. We validate HealthQ using two custom datasets constructed from public medical datasets, ChatDoctor and MTS-Dialog, and demonstrate its robustness across multiple LLM judge models, including GPT-3.5, GPT-4, and Claude. Our contributions are threefold: we present the first systematic framework for assessing questioning capabilities in healthcare conversations, establish a model-agnostic evaluation methodology, and provide empirical evidence linking high-quality questions to improved patient information elicitation.
I Introduction
@@
The integration of Artificial Intelligence (AI) in healthcare has transformed personal healthcare, enabling advancements in diagnostic accuracy [1], treatment personalization [2, 3], and patient management [4, 5, 6, 7]. At the core of this transformation are Large Language Models (LLMs), which exhibit exceptional abilities in understanding and generating human-like text. These capabilities have been harnessed in healthcare applications such as virtual assistants, automated diagnostic tools, and clinical reasoning systems [8]. Much of the existing research has focused on enhancing LLM-based question-answering (QA) capabilities, which are pivotal for delivering timely and accurate responses to patient inquiries. Datasets and benchmarks such as MedQA [7], emrQA [9], and MedMCQA [10] have been developed to train and evaluate the QA abilities of LLMs in healthcare.
Recent advancements have demonstrated that LLMs can be organized into structured workflows, referred to as LLM healthcare chains, to address complex interactions [8, 11, 12, 13], These chains represent a structured sequence of operations designed to manage multi-step tasks in medical contexts, integrating LLMs with additional tools, external knowledge bases, and reasoning mechanisms. Unlike standalone models, LLM healthcare chains leverage components such as Retrieval-Augmented Generation (RAG)[14] for retrieving domain-specific information, Chain of Thought (CoT) reasoning[15] for step-by-step logical processing, and reflective chains for iterative refinement of outputs. In the context of healthcare, LLM healthcare chains encompass workflows tailored to retrieve relevant medical knowledge, interpret patient inputs, and generate diagnostic or follow-up questions. For example, when a patient presents symptoms, the chain might retrieve pertinent medical data, analyze patient responses to identify gaps, and generate contextually appropriate questions to gather additional information. These chains facilitate dynamic, multi-turn interactions, ensuring that conversations are not only relevant but also contextually rich, making them particularly well-suited for applications in diagnostic reasoning, patient education, and treatment planning. This structured and adaptive approach distinguishes LLM healthcare chains from static single-turn question-answering systems, providing a robust framework for managing complex healthcare conversations.
Despite the progress in QA systems, effective patient care demands more than the ability to answer questions. Healthcare interactions often require LLM healthcare chains to proactively ask pertinent and context-aware questions, eliciting critical information such as symptoms, medical history, and lifestyle factors. This questioning capability is crucial for accurate diagnosis and personalized treatment, yet it remains significantly under-explored [16, 17]. Existing benchmarks, which primarily evaluate static, single-turn QA performance, are insufficient for assessing the interactive and information-seeking abilities of LLMs. The MediQ [11] benchmark highlights the importance of transitioning from static to interactive paradigms, where LLMs must actively gather missing details via follow-up questions to achieve reliable clinical reasoning. These insights motivate the need for a dedicated framework to evaluate the question-asking capabilities of LLM healthcare chains in realistic settings.
To address this gap, we propose HealthQ, a novel framework for evaluating the questioning capabilities of LLM healthcare chains. HealthQ assesses both the quality of questions generated during interactions and their effectiveness in eliciting informative responses. Our framework employs a multi-dimensional evaluation methodology, incorporating traditional Natural Language Processing (NLP) metrics (e.g., ROUGE and Named Entity Recognition), as well as LLM-based interrogation metrics (e.g., relevance, specificity, and specificity). We validate the framework using custom datasets derived from public medical note datasets such as ChatDoctor and MTS-Dialog, ensuring realistic and comprehensive evaluation scenarios.
Our contributions are threefold:
@@
    We conduct the first comprehensive study on the questioning capabilities of LLM healthcare chains, validating performance across multiple LLMs (GPT-3.5, GPT-4, Claude) and establishing metrics for evaluating question quality in healthcare contexts.
@@
    We develop a dataset generation pipeline with validated parsing processes, enabling systematic evaluation of question-asking capabilities in interactive patient interactions.
@@
    We propose an evaluation framework that integrates multiple LLM chain architectures and NLP metrics, demonstrating consistent performance across different LLM judges and providing reproducible benchmarks for medical question generation.
The code, dataset generation pipeline, and evaluation framework are publicly available at https://github.com/wangziyu99/HealthQ-LLM-Healthcare-Benchmark.
II Related Work
II-A Digital Healthcare Interventions
The advent of digital healthcare interventions has revolutionized patient care by leveraging technologies such as wearable devices [18, 19, 20, 21, 22], and AI-driven analytics [8, 23]. These interventions enable remote monitoring, personalized treatment plans, and predictive analytics, improving patient outcomes and reducing healthcare costs [24, 25]. Novel AI integration, including machine learning and natural language processing, enhances diagnostic accuracy and patient management [26, 4]. Machine learning algorithms predict disease progression and personalize treatments, while natural language processing interprets unstructured clinical data, enabling efficient information retrieval and decision support [27]. These advancements highlight the potential of digital healthcare to improve the efficiency and effectiveness of care delivery.
II-B LLMs in Healthcare
LLMs such as GPT-4 have shown remarkable proficiency in generating human-like text and understanding complex language constructs [28, 29, 30]. These capabilities have been harnessed in various healthcare applications, including automated medical record summarization, virtual health assistants, and clinical decision support systems [31, 32, 33]. Automated medical record summarization condenses patient records into concise summaries, saving time for healthcare providers. Virtual health assistants interact with patients using natural language, providing answers to medical queries and reminders for medication, thereby enhancing patient engagement. Clinical decision support systems leverage the extensive knowledge embedded in LLMs to assist in diagnosing diseases and suggesting treatment options. LLMs streamline clinical workflows and improve patient engagement by providing accurate and timely responses to medical queries, ultimately enhancing the overall quality of care.
II-C Interactive Questioning with LLM Healthcare Chains
@@
Several approaches have been introduced to refine the questioning capabilities of LLMs. Techniques such as Chain of Thought (CoT) [15] prompting, reflective chains [35], and RAG [14] integrate reasoning processes that help generate more contextually relevant questions. CoT prompting allows LLMs to break down complex queries into intermediate steps, leading to more precise and targeted questioning. Reflective chains enable the model to iteratively improve the questions by reflecting on previous intermediate answers, thus enhancing the relevance of subsequent queries. RAG, on the other hand, combines retrieval and generation processes to ensure that questions are informed by the most relevant information retrieved from a knowledge base. Despite these advancements, most existing work has primarily focused on optimizing task completion and response accuracy, with less attention given to the development of sophisticated questioning strategies that can drive more informative patient interactions. Our work addresses this gap by specifically evaluating the questioning capabilities of LLM healthcare chains, contributing a novel framework that benchmarks their ability to engage in meaningful and effective patient dialogues.
III Method
III-A Overview of HealthQ
@@
III-B Data
HealthQ necessitates a general data processing mechanism due to the unstructured nature of medical notes and the often ambiguous doctor-patient conversations. Our data reformatting serves two critical purposes: (1) establishing ground truth for evaluating generated questions by identifying what information doctors sought in successful consultations, and (2) creating separate knowledge bases for virtual patient simulation and question evaluation. When doctors ask effective questions in real consultations, they uncover specific patient information - this relationship between questions and discovered information forms our evaluation basis.
To accurately evaluate the questioning capabilities of LLM healthcare chains, we extract two key components from each consultation: the patient’s known information (symptoms, feelings, and history) and the doctor’s information-seeking behavior (what information the doctor aimed to uncover through their questions). This paired format allows us to assess whether generated questions would effectively uncover similar critical information. For example, if the original consultation shows a doctor successfully uncovering fever details through specific questions, we can evaluate whether our LLM chains generate questions that would similarly reveal this important information. To illustrate this relationship between original consultations and our evaluation framework, Table I shows an example of the original and processed data formats.
III-B1 Dataset
@@
TABLE I: Example of original and processed data. The original format includes detailed patient background, symptoms, diagnosis, history, plan of action, and dialogue. The processed format extracts essential patient known knowledge and reformulates doctor question statements for clarity and consistency.
@@
III-B2 Data Processing
As depicted in Figure 1, we partition the medical notes into two sets: one for constructing the knowledge base and another for evaluation. The first set is indexed in a vector database (VDB), serving as a reference knowledge base that LLM chains can query during operation - similar to how RAG systems typically operate. This is not traditional model training, but rather creation of a searchable medical knowledge repository. The evaluation set remains separate to assess the chains’ questioning capabilities on entirely new patient scenarios.
Specifically, we construct the knowledge base using FAISS [^1] https://github.com/facebookresearch/faiss for efficient similarity search. The customized collate functions specify comparison fields and utilize metadata for enhanced information retrieval. This flexibility accommodates diverse datasets with varying forms and topics. During operation, multiple VDB searching functions can be provided as tools without merging, allowing the LLM to select the appropriate knowledge source. VDBs work by finding the nearest vectors, which are represented by embedding models. Each vector database entry is defined as:
@@ III-B2 Data Processing
Each entry in the vector database, referred to as a record, is structured to include both the content and the associated metadata. The collate function processes the input data to create these records:
$\text{Record}_{i}=(\text{Content}_{i},\text{Metadata}_{i})=\text{Collate}(\text{Data}_{\text{kb}}(i))$
This knowledge base facilitates efficient retrieval and processing of medical information during question generation. For the evaluation set, we employed Claude [^2] https://claude.ai/ to parse patient notes into structured formats, producing pairs of patient known knowledge and doctor question statements. The final evaluation ground truth, $G$, is thus composed of:
$GT=\{\text{LLM\_Parser}(\text{Data}_{\text{eval}}(i))\}$
III-C Baseline Healthcare LLM Chains
Healthcare LLM chains are structured workflows that utilize LLMs to simulate patient-doctor interactions by generating diagnostic questions and interpreting responses. These chains serve as frameworks for improving the relevance, specificity, and informativeness of questions in healthcare conversations, enabling precise information gathering. In this work, we implement several state-of-the-art LLM chains, each designed with unique mechanisms for question generation.
@@ III-C Baseline Healthcare LLM Chains
TABLE II: LLM Judge Interrogation Metrics (Scores range from 0 to 1)
Metric Description
Specificity Evaluates how precisely the question addresses the patient’s described symptoms and medical context, ensuring the question targets clinically relevant details.
Usefulness Assesses the diagnostic value of the question, focusing on its potential to elicit new or critical information that contributes to a better understanding of the patient’s condition.
Relevance Determines the alignment of the question with the patient’s symptoms, medical history, and overall context, ensuring it stays pertinent to the clinical scenario.
Coverage Measures how comprehensively the question explores the available patient information, ensuring that no critical aspects of the patient’s condition are overlooked.
Fluency Evaluates the clarity and coherence of the question’s formulation, ensuring it is expressed in clear and natural medical language appropriate for healthcare communication.
III-D Evaluation Framework
@@
  • LLM Judge Interrogation: This metric evaluates the questions generated by the LLM healthcare chains based on specificity, usefulness, relevance, coverage, and fluency, as shown in Table II. An LLM judge $J$ assesses each question $Q$ using the patient knowledge $K$, providing scores according to these criteria. The interrogation metrics are calculated as follows:
@@
    Evaluate the Question ($Q$) using Judge ($J(Q,K)$): Evaluate the generated question using an LLM judge based on criteria such as relevance, specificity, and informativeness, as detailed in Table II.